# Conformally invariant boundary arcs in double dimers

### v1

@@ 2.2. Proof of Kenyon’s formula
×$\quad\times\prod_{\mathrm{non~{}trivial~{}arcs}\ A}(-1)^{r_{1}+s_{1}+\cdots+r_{k}+s_{k}}(-1)^{\mathbbm{1}\{b_{A}>w_{A}\}},$
@@
Let $m$ be a dimer configuration on $G^{\times}$, and let $D_{m}$ be the intersecting curves given by $m$ as described above. The vertices of $G^{\times}$ are arranged clockwise on the boundary of a disc $D$ according to the order $<$, and each pair of vertices in an edge of $m$ are joined by a curve in $D_{m}$. Recall $k(m)$ is the number of crossings of these curves. One can permute the pairs of bulk vertices $(v,1),(v,2)$ with each other in $D_{m}$ and preserve $(-1)^{k(m)}$. For example, if the vertices $(v,1),(v,2),(u,1),(u,2)$ appear consecutively on the boundary of $D_{m}$, permuting them so they appear in the order $(u,1),(u,2),(v,1),(v,2)$ preserves $(-1)^{k(m)}$. We will use this property several times in the rest of the proof. The curves connecting these vertices now only cross each other and no other curve in $D_{m}$, as the vertices are consecutive.
@@
Figure 8. The vertices and edges of $m_{0}$ arranged in some part of $D_{m_{0}}$; the ordering of the vertices here is given by $<$.
and one can see that there are exactly $12k-$ 1 crossings, hence $k_{m_{0},C}=-1$, and (5) holds. Note that in Figure 8 the vertices are ordered in $D_{m_{0}}$ according to $<$, whereas in Figure 7 they are arranged as in $G^{\times}$.
@@
The case drawn is the case where $w_{A}<b_{A}$, and there are $12k-$ 1 crossings. In the case $w_{A}>b_{A}$, the positions of $w_{A}$ and $b_{A}$ are exchanged, and there is exactly one more crossing. Hence (6) holds. By the same inductive argument as used for loops above, (6) holds for all dimer configurations $m$ and all arcs $A$ in $\omega=p(m)$.
@@
$\begin{split}\mathrm{Pf}K&=\sum_{m}(-1)^{k(m)}\prod_{\begin{subarray}{c}xy\in m\\ x<y\end{subarray}}K_{x,y}\\ &=\sum_{\omega\in\Omega}\sum_{m\in p^{-1}(\omega)}(-1)^{k(m)}\prod_{\mathrm{doubled~{}edges}\ e}\prod_{\begin{subarray}{c}xy\in e\\ x<y\end{subarray}}K_{x,y}\prod_{\mathrm{loops}\ C}\prod_{\begin{subarray}{c}xy\in C\\ x<y\end{subarray}}K_{x,y}\prod_{\mathrm{arcs}\ A}\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}K_{x,y}\\ &=\sum_{\omega\in\Omega}\sum_{\mathrm{indices}}\prod_{\mathrm{doubled~{}edge}\ e}(-1)^{i_{e}}\prod_{\begin{subarray}{c}xy\in e\\ x<y\end{subarray}}K_{x,y}\prod_{\mathrm{loops}\ C}(-1)^{1+i_{1}+\cdots+j_{k}}\prod_{\begin{subarray}{c}xy\in C\\ x<y\end{subarray}}K_{x,y}\\ &\quad\prod_{\mathrm{trivial~{}arcs}\ A}\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}K_{x,y}\prod_{\begin{subarray}{c}\mathrm{non-trivial}\\ \mathrm{arcs}\ A\end{subarray}}(-1)^{r_{1}+s_{1}+\cdots+r_{k}+s_{k}+\mathbbm{1}\{b_{A}>w_{A}\}}\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}K_{x,y},\end{split}$ (7)
@@
$\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}K_{x,y}=(-1)^{\mathbbm{1}\{w_{A}>b_{A}\}}\xi_{w_{A},b_{A}}\nu_{w_{A},b_{A}}=\nu_{w_{A},b_{A}}=1,$
@@
We now use that for matrices $\phi\in\mathrm{SL}_{2}(\mathbb{C})$, $\phi_{3-i,3-j}=(-1)^{i+j}\phi^{-1}_{j,i}$, which one can verify by hand. We replace each factor $(\phi_{u_{l+1},v_{l}})_{3-i_{l+1},3-j_{l}}$ with $(-1)^{j_{l}+i_{l+1}}(\phi_{v_{l},u_{l+1}})_{j_{l},i_{l+1}}$, and the above becomes:
@@
Now the factor $(-1)^{i_{1}+\cdots+j_{k}}$ in the above cancels with the same factor on the left hand side of (8), and the sum over all $i_{1},j_{1},\dots,i_{k},j_{k}$ produces the trace. The product of factors $\xi_{e}$ around a loop $C$ is $(-1)^{\tfrac{|C|}{2}+1}$ by the Kasteleyn phases; the $+1$ in this exponent cancels with the remaining factor $-1$ in the left hand side of (8). Hence (8) holds.
@@
$\sum_{r_{1},s_{1},\dots,r_{k},s_{k}}(-1)^{r_{1}+s_{1}+\cdots+r_{k}+s_{k}+\mathbbm{1}\{b_{A}>w_{A}\}}\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}K_{x,y}=(-1)^{\tfrac{1}{2}(|A|-1)}\psi_{w_{A}}^{\intercal}\phi_{A}\psi_{b_{A}},$ (9)
@@
One then only needs that $\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}\xi_{x,y}=(-1)^{\mathbbm{1}\{b_{A}>w_{A}\}}(-1)^{\tfrac{1}{2}(|A|-1)}$. Recall that we constructed $\xi$ to be a set of Kasteleyn phases such that $\xi_{w,b}=(-1)^{\mathbbm{1}\{w>b\}}$ on boundary edges $\{w,b\}$ (apart perhaps from some edge $e_{0}$). Let $C_{A}$ be the loop in $G$ made up of an arc from $w_{A}$ to $b_{A}$, and then the boundary edges from $w_{A}$ to $b_{A}$ (not in the direction including $e_{0}$), which we denote $\partial C_{A}$. One can show from the above that $\prod_{\begin{subarray}{c}xy\in\partial C_{A}\\ x<y\end{subarray}}\xi_{x,y}=(-1)^{\mathbbm{1}\{b_{A}>w_{A}\}}(-1)^{\tfrac{1}{2}(|\partial C_{A}|+1)}$. Then
$\begin{split}\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}\xi_{x,y}&=\prod_{\begin{subarray}{c}xy\in C_{A}\\ x<y\end{subarray}}\xi_{x,y}\prod_{\begin{subarray}{c}xy\in\partial C_{A}\\ x<y\end{subarray}}\xi_{x,y}\\ &=(-1)^{\tfrac{1}{2}|C_{A}|+1}(-1)^{\mathbbm{1}\{b_{A}>w_{A}\}}(-1)^{\tfrac{1}{2}(|\partial C_{A}|+1)}\\ &=(-1)^{\mathbbm{1}\{b_{A}>w_{A}\}}(-1)^{\tfrac{1}{2}(|A|-1)}.\end{split}$
@@ 3.1.1. Temperleyan and piecewise Temperleyan approximations
For every $\varepsilon>0$, $U_{\varepsilon}$ is a graph with vertex set $V(U_{\varepsilon})\subset\bar{U}\cap\varepsilon\mathbb{Z}^{2}$ and edges connecting points at distance $\epsilon$: if $u,v\in V(U_{\varepsilon})$, $u\sim v$ in $U_{\varepsilon}$ if and only if $u\sim v$ in $\varepsilon\mathbb{Z}^{2}$. For every $\varepsilon>0$, $U_{\varepsilon}$ is simply connected: if we cover every vertex of $U_{\varepsilon}$ by a square of side $\varepsilon$ centered at the vertex, we obtain a simply connected set. The vertex boundary of $U_{\varepsilon}$ (i.e. the vertices of $U_{\varepsilon}$ which are connected in $\varepsilon\mathbb{Z}^{2}$ to at least one vertex not in $U_{\varepsilon}$) is within $O(\varepsilon)$ of $\partial U$.
@@
From now on, let $U^{r}$ be a simply connected open set with smooth boundary which is symmetric by reflection along the horizontal axis. For such symmetric sets $U^{r}$, we denote by $U=U\cap\mathbb{R}\times\mathbb{R}_{>0}$ the restriction of $U^{r}$ to the strict upper half plane. Let $z\in U$ be fixed. For technical reasons, we assume that there exists $z_{\partial}\in\partial U^{r}$ and $\delta>0$ such that the boundary of $U^{r}$ is horizontal in the neighbourhood $B(z_{\partial},\delta)$. We also assume that $U^{r}$ is contained in the half-plane below this flat horizontal boundary.
@@
For every $\varepsilon>0$, let $U_{\varepsilon}$ be the graphs obtained by removing from $U_{\varepsilon}^{r}$ the vertices in the strict lower half plane and the edges incident to them. The sequence $(\mathcal{G}_{\varepsilon})_{\varepsilon>0}=(\mathcal{G}_{\varepsilon}^{1})_{\varepsilon>0}$ obtained by restricting $\mathcal{G}_{\varepsilon}^{r}$ to the upper half plane (removing all vertices of $\mathcal{G}_{\varepsilon}^{r}$ in ×$\mathbb{R}\times\mathbb{R}_{<0}$, and the edges incident to them) is a Temperleyan approximation of the open set $U$. Dimer configurations of $\mathcal{G}_{\varepsilon}=\mathcal{G}_{\varepsilon}^{1}$ are in bijection with spanning trees of $U_{\varepsilon}$ rooted at $b_{0}$.
@@
The sequence $(\mathcal{G}_{\varepsilon}^{2})_{\varepsilon>0}$ obtained by restricting $\mathcal{G}_{\varepsilon}^{r}$ to the strict upper half plane (i.e. removing all vertices of $\mathcal{G}_{\varepsilon}^{r}$ in ×$\mathbb{R}\times\mathbb{R}_{\leq 0}$ and the edges incident to them) is a piecewise Temperleyan approximation of the open set $U$. Indeed, it has exactly two convex white corner, $v_{l}^{*}$ and $v_{r}^{*}$ respectively on the left and right of the horizontal axis, see the definition in Section 5.1 of [20].
@@ 3.1.2. Loops and arcs on a Temperleyan approximation
3.1.3. Discrete and continuous Green’s function.
In this paragraph, we recall the notation and results of [17] on continuous Green’s functions. Let $g:u,v\in U^{r}\to g(u,v)\in\mathbb{R}$ be the continuous Green’s function with Dirichlet boundary conditions on $U^{r}$. Let $\tilde{g}:u,v\in U^{r}\to g(u,v)\in\mathbb{C}$ be the analytic function of $v$ whose real part is the Dirichlet Green’s function. Define its Wirtinger derivatives
@@
Recall that the boundary of $U^{r}$ is horizontal in the ball $B(z_{\partial},\delta/2)$ and that $U^{r}$ is contained in the half-plane below this horizontal boundary. Denote by $s$ the reflection along this horizontal boundary. Since the Green’s function $g$ has Dirichlet boundary conditions, i.e. it is 00 on the boundary of $U^{r}$, the Schwarz reflection principle (see for example Section 6.5 of [1]) implies that $i\tilde{g}(u,v)$ as a function of $v$ extends to an analytic (except at $u$ and $s(u)$) function on $U^{r}\cup s(U^{r})$. Hence $F_{+}$ and $F_{-}$ also extend analytically (as functions of $v$) to this domain, in particular they are well-defined at $z_{\partial}$ and they remain bounded near $z_{\partial}$ (and so do their derivatives). In the proof of Lemma B.3, we will detail a discrete version of this argument developed by Kenyon in the proof of Theorem 14 of [15].
@@
∫⋯∫$\idotsint_{\gamma}\prod_{i=0}^{n-1}F^{(\sigma_{i})}_{-\sigma_{i}\sigma_{i+1}}(\overline{z_{i+1}},z_{i})\prod_{i=0}^{n-1}dz^{(\sigma_{i})}_{i},∫⋯$ (11)
where $dz_{k_{i}}^{(1)}=dz_{k_{i}}$ and $dz_{k_{i}}^{(-1)}=d\overline{z_{k_{i}}}$, and similar for $F_{+}$ and $F_{-}$, and where the subscript $-\sigma_{i}\sigma_{i+1}$ should be read as $\pm$ when $-\sigma_{i}\sigma_{i+1}=\pm 1$, respectively. This is well defined since the integral of an analytic function does not depend on the path.
@@
[…]
Note that $c_{n}(z,z_{\partial})$ is always real, since the terms in the sum corresponding to $\sigma$ and $-\sigma$ are complex conjugates. Moreover, since $F_{+}$ and $F_{-}$ are analytic on $U$ and remain bounded near $z_{\partial}$, the $c_{n}(z,z_{\partial})$ grow at most exponentially: there exists an absolute constant $C$ (depending only on $\gamma$ and $U$) such that
$\forall n\in\mathbb{Z}_{\geq 0},~{}c_{n}(z,z_{\partial})\leq C^{n}.$ (13)
@@
$\mathbbm{E}_{\varepsilon}\left[\binom{\frac{n_{\varepsilon}(z)-o_{\varepsilon}(z)}{2}}{n}\right]\geq\mathbbm{E}_{\varepsilon}\left[\binom{\frac{n_{\varepsilon}(z)-1}{2}}{n}\right]$ $\geq\mathbbm{E}_{\varepsilon}\left[\mathbbm{1}_{n_{\varepsilon}(z)\geq 2n}\frac{(n_{\varepsilon}(z)-2n)^{n}}{2^{n}n!}\right]$
$\geq\mathbbm{E}_{\varepsilon}\left[\mathbbm{1}_{n_{\varepsilon}(z)\geq 4n}\frac{n_{\varepsilon}(z)^{n}}{4^{n}n!}\right]\geq\frac{1}{4^{n}n!}(\mathbbm{E}_{\varepsilon}[n_{\varepsilon}(z)^{n}]-(4n)^{n}).$
@@
since there are at most $n$ choices for $j_{1}$, $n/2$ choices for $j_{2}$ etc. Using Stierling’s formula, we conclude that $m_{n}\leq C^{n}n^{n}$ for another absolute constant $C$, hence Carleman’s condition is verified and $n_{\varepsilon}(z)$ converges towards a random variable $n(z)$ which is uniquely determined by its moments $m_{n}$. Since this moments are conformally invariant (they are expressed as integrals of conformal quantities), the law of $n(z)$ is also conformally invariant. ∎
In the particular case $U^{r}=[-\infty,+\infty]\times[-\pi/2,\pi/2]$ where the Green’s function has a simple explicit expression, the computations turn out to be particularly simple.
@@
The infinite strip does not fill the technical hypothesis of our theorem since it is not a bounded open set, although it fills the other conditions: it is symmetric and has a horizontal boundary. Nonetheless, since the formula for the moments depends only on conformally invariant quantities, the moments can be computed in any simply connected open set, and the computations transfered to any other domain by a Riemann map. Besides, if we take $U_{\varepsilon}^{r}=\varepsilon\mathcal{Z}^{2}\cap[-1/\varepsilon,1/\varepsilon]\times[-\pi/2,\pi/2]$, the proof of Theorem 3.4 still holds, at least for the folded dimer model: the only part of the proof that has to be modified is Lemma B.3 of the Appendix. The proof of this lemma still works because, to show that $K^{-1}(b,w)$ is close to the partial derivative of the Green’s function $g$, we only need that $g_{\mathbb{R}^{2}}$ (the full-plane Green’s function) or $g_{\mathbb{H}}$ (the half-plane Green’s function) and $g$ are $O(\varepsilon)$ on the boundary of $\mathcal{G}_{\varepsilon}^{r}$ to apply the Harnack lemma. This is also true here.
@@
[…]

### v2

@@ 2.2. Proof of Kenyon’s formula
$\quad\times\prod_{\mathrm{non~trivial~arcs}\ A}(-1)^{r_{1}+s_{1}+\cdots+r_{k}+s_{k}}(-1)^{\mathds{1}\{b_{A}>w_{A}\}},$
@@
Let $m$ be a dimer configuration on $G^{\times}$, and let $D_{m}$ be the intersecting curves given by $m$ as described above. The vertices of $G^{\times}$ are arranged clockwise on the boundary of a disc $D$ according to the order <<, and each pair of vertices in an edge of $m$ are joined by a curve in $D_{m}$. Recall $k(m)$ is the number of crossings of these curves. One can permute the pairs of bulk vertices $(v,1),(v,2)$ with each other in $D_{m}$ and preserve $(-1)^{k(m)}$. For example, if the vertices $(v,1),(v,2),(u,1),(u,2)$ appear consecutively on the boundary of $D_{m}$, permuting them so they appear in the order $(u,1),(u,2),(v,1),(v,2)$ preserves $(-1)^{k(m)}$. We will use this property several times in the rest of the proof. The curves connecting these vertices now only cross each other and no other curve in $D_{m}$, as the vertices are consecutive.
@@
Figure 9. The vertices and edges of $m_{0}$ arranged in some part of $D_{m_{0}}$; the ordering of the vertices here is given by <<.
and one can see that there are exactly 2k12k-1 crossings, hence $k_{m_{0},C}=-1$, and (5) holds. Note that in Figure 9 the vertices are ordered in $D_{m_{0}}$ according to <<, whereas in Figure 8 they are arranged as in $G^{\times}$.
@@
The case drawn is the case where $w_{A}<b_{A}$, and there are 2k12k-1 crossings. In the case $w_{A}>b_{A}$, the positions of $w_{A}$ and $b_{A}$ are exchanged, and there is exactly one more crossing. Hence (6) holds. By the same inductive argument as used for loops above, (6) holds for all dimer configurations $m$ and all arcs $A$ in $\omega=p(m)$.
@@
$\begin{split}\mathrm{Pf}K&=\sum_{m}(-1)^{k(m)}\prod_{\begin{subarray}{c}xy\in m\\ x<y\end{subarray}}K_{x,y}\\ &=\sum_{\omega\in\Omega}\sum_{m\in p^{-1}(\omega)}(-1)^{k(m)}\prod_{\mathrm{doubled~edges}\ e}\prod_{\begin{subarray}{c}xy\in e\\ x<y\end{subarray}}K_{x,y}\prod_{\mathrm{loops}\ C}\prod_{\begin{subarray}{c}xy\in C\\ x<y\end{subarray}}K_{x,y}\prod_{\mathrm{arcs}\ A}\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}K_{x,y}\\ &=\sum_{\omega\in\Omega}\sum_{\mathrm{indices}}\prod_{\mathrm{doubled~edge}\ e}(-1)^{i_{e}}\prod_{\begin{subarray}{c}xy\in e\\ x<y\end{subarray}}K_{x,y}\prod_{\mathrm{loops}\ C}(-1)^{1+i_{1}+\cdots+j_{k}}\prod_{\begin{subarray}{c}xy\in C\\ x<y\end{subarray}}K_{x,y}\\ &\quad\prod_{\mathrm{trivial~arcs}\ A}\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}K_{x,y}\prod_{\begin{subarray}{c}\mathrm{non-trivial}\\ \mathrm{arcs}\ A\end{subarray}}(-1)^{r_{1}+s_{1}+\cdots+r_{k}+s_{k}+\mathds{1}\{b_{A}>w_{A}\}}\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}K_{x,y},\end{split}$ (7)
@@
$\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}K_{x,y}=(-1)^{\mathds{1}\{w_{A}>b_{A}\}}\xi_{w_{A},b_{A}}\nu_{w_{A},b_{A}}=\nu_{w_{A},b_{A}}=1,$
@@
We now use that for matrices $\phi\in\mathrm{SL}_{2}(\mathbb{C})$, $\phi_{3-i,3-j}=(-1)^{i+j}(\phi^{-1})_{j,i}$, which one can verify by hand. We replace each factor $(\phi_{u_{l+1},v_{l}})_{3-i_{l+1},3-j_{l}}$ with $(-1)^{j_{l}+i_{l+1}}(\phi_{v_{l},u_{l+1}})_{j_{l},i_{l+1}}$, and the above becomes:
@@
Now the factor $(-1)^{i_{1}+\cdots+j_{k}}$ in the above cancels with the same factor on the left hand side of (8), and the sum over all $i_{1},j_{1},\dots,i_{k},j_{k}$ produces the trace. The product of factors $\xi_{e}$ around a loop $C$ is $(-1)^{\tfrac{|C|}{2}+1}$ by the Kasteleyn phases; the +1+1 in this exponent cancels with the remaining factor 1-1 in the left hand side of (8). Hence (8) holds.
@@
$\sum_{r_{1},s_{1},\dots,r_{k},s_{k}}(-1)^{r_{1}+s_{1}+\cdots+r_{k}+s_{k}+\mathds{1}\{b_{A}>w_{A}\}}\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}K_{x,y}=(-1)^{\tfrac{1}{2}(|A|-1)}\psi_{w_{A}}^{\intercal}\phi_{A}\psi_{b_{A}},$ (9)
@@
One then only needs that $\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}\xi_{x,y}=(-1)^{\mathds{1}\{b_{A}>w_{A}\}}(-1)^{\tfrac{1}{2}(|A|-1)}$. Recall that we constructed $\xi$ to be a set of Kasteleyn phases such that $\xi_{w,b}=(-1)^{\mathds{1}\{w>b\}}$ on boundary edges $\{w,b\}$ (apart perhaps from some edge $e_{0}$). Let $C_{A}$ be the loop in $G$ made up of an arc from $w_{A}$ to $b_{A}$, and then the boundary edges from $w_{A}$ to $b_{A}$ (not in the direction including $e_{0}$), which we denote $\partial C_{A}$. One can show from the above that $\prod_{\begin{subarray}{c}xy\in\partial C_{A}\\ x<y\end{subarray}}\xi_{x,y}=(-1)^{\mathds{1}\{b_{A}>w_{A}\}}(-1)^{\tfrac{1}{2}(|\partial C_{A}|+1)}$. Then
$\begin{split}\prod_{\begin{subarray}{c}xy\in A\\ x<y\end{subarray}}\xi_{x,y}&=\prod_{\begin{subarray}{c}xy\in C_{A}\\ x<y\end{subarray}}\xi_{x,y}\prod_{\begin{subarray}{c}xy\in\partial C_{A}\\ x<y\end{subarray}}\xi_{x,y}\\ &=(-1)^{\tfrac{1}{2}|C_{A}|+1}(-1)^{\mathds{1}\{b_{A}>w_{A}\}}(-1)^{\tfrac{1}{2}(|\partial C_{A}|+1)}\\ &=(-1)^{\mathds{1}\{b_{A}>w_{A}\}}(-1)^{\tfrac{1}{2}(|A|-1)}.\end{split}$
@@ 3.1.1. Temperleyan and piecewise Temperleyan approximations
For every $\varepsilon>0$, $U_{\varepsilon}$ is a graph with vertex set $V(U_{\varepsilon})\subset\bar{U}\cap\varepsilon\mathbb{Z}^{2}$ and edges connecting points at distance $\varepsilon$: if $\in V(U_{\varepsilon}), uvu\sim v$ in $U_{\varepsilon}$ if and only if $u\sim v$ in $\varepsilon\mathbb{Z}^{2}$. For every $\varepsilon>0$, $U_{\varepsilon}$ is simply connected: if we cover every vertex of $U_{\varepsilon}$ by a square of side $\varepsilon$ centered at the vertex, we obtain a simply connected set. The vertex boundary of $U_{\varepsilon}$ (i.e. the vertices of $U_{\varepsilon}$ which are connected in $\varepsilon\mathbb{Z}^{2}$ to at least one vertex not in $U_{\varepsilon}$) is within $O(\varepsilon)$ of $\partial U$.
@@
From now on, let $U^{r}$ be a simply connected open set with smooth boundary which is symmetric by reflection along the horizontal axis. For such symmetric sets $U^{r}$, we denote by $U=U\cap(\mathbb{R}\times\mathbb{R}_{>0})$ the restriction of $U^{r}$ to the strict upper half plane. Let $z\in U$ be fixed. For technical reasons, we assume that there exists $z_{\partial}\in\partial U^{r}$ and $\delta>0$ such that the boundary of $U^{r}$ is horizontal in the neighbourhood $B(z_{\partial},\delta)$. We also assume that $U^{r}$ is contained in the half-plane below this flat horizontal boundary.
@@
For every $\varepsilon>0$, let $U_{\varepsilon}$ be the graphs obtained by removing from $U_{\varepsilon}^{r}$ the vertices in the strict lower half plane and the edges incident to them. The sequence $(\mathcal{G}_{\varepsilon})_{\varepsilon>0}=(\mathcal{G}_{\varepsilon}^{1})_{\varepsilon>0}$ obtained by restricting $\mathcal{G}_{\varepsilon}^{r}$ to the upper half plane (removing all vertices of $\mathcal{G}_{\varepsilon}^{r}$ in $\mathbb{R}\times\mathbb{R}_{<0}$, and the edges incident to them) is a Temperleyan approximation of the open set $U$. Dimer configurations of $\mathcal{G}_{\varepsilon}=\mathcal{G}_{\varepsilon}^{1}$ are in bijection with spanning trees of $U_{\varepsilon}$ rooted at $b_{0}$.
@@
The sequence $(\mathcal{G}_{\varepsilon}^{2})_{\varepsilon>0}$ obtained by restricting $\mathcal{G}_{\varepsilon}^{r}$ to the strict upper half plane (i.e. removing all vertices of $\mathcal{G}_{\varepsilon}^{r}$ in $\mathbb{R}\times\mathbb{R}_{\leq 0}$ and the edges incident to them) is a piecewise Temperleyan approximation of the open set $U$. Indeed, it has exactly two convex white corner, $v_{l}^{*}$ and $v_{r}^{*}$ respectively on the left and right of the horizontal axis, see the definition in Section 5.1 of [19].
@@ 3.1.2. Loops and arcs on a Temperleyan approximation
3.1.3. Discrete and continuous Green function.
In this paragraph, we recall the notation and results of [16] on continuous Green functions. Let $g:u,v\in U^{r}\to g(u,v)\in\mathbb{R}$ be the continuous Green function with Dirichlet boundary conditions on $U^{r}$. Let $\tilde{g}:u,v\in U^{r}\to g(u,v)\in\mathbb{C}$ be the analytic function of $v$ whose real part is the Dirichlet Green function. Define its Wirtinger derivatives
@@
Recall that the boundary of $U^{r}$ is horizontal in the ball $B(z_{\partial},\delta/2)$ and that $U^{r}$ is contained in the half-plane below this horizontal boundary. Denote by $s$ the reflection along this horizontal boundary. Since the Green function $g$ has Dirichlet boundary conditions, i.e. it is 0 on the boundary of $U^{r}$, the Schwarz reflection principle (see for example Section 6.5 of [1]) implies that $i\tilde{g}(u,v)$ as a function of $v$ extends to an analytic (except at $u$ and $s(u)$) function on $U^{r}\cup s(U^{r})$. Hence $F_{+}$ and $F_{-}$ also extend analytically (as functions of $v$) to this domain, in particular they are well-defined at $z_{\partial}$ and they remain bounded near $z_{\partial}$ (and so do their derivatives). In the proof of Lemma B.3, we will detail a discrete version of this argument developed by Kenyon in the proof of Theorem 14 of [14].
@@
$\idotsint_{\gamma}\prod_{i=0}^{n-1}F^{(\sigma_{i})}_{-\sigma_{i}\sigma_{i+1}}(\overline{z_{i+1}},z_{i})\prod_{i=0}^{n-1}dz^{(\sigma_{i})}_{i}$, (11)
where $dz_{k_{i}}^{(1)}=dz_{k_{i}}$ and $dz_{k_{i}}^{(-1)}=d\overline{z_{k_{i}}}$, and similar for $F_{+}$ and $F_{-}$, and where the subscript $\sigma_{i}\sigma_{i+1}$ should be read as $\pm$ when $\sigma_{i}\sigma_{i+1}=\pm 1$, respectively. This is well defined since the integral of an analytic function does not depend on the path.
@@
$\left\{\begin{array}[]{ll}c_{1}(z,z_{\partial})&=-4\Im\left\{\int_{\gamma}F_{-}(\overline{z},z)dz\right\}\\ c_{2}(z,z_{\partial})&=4\Re\left\{\int\int_{\gamma}F_{+}(\overline{z_{1}},z_{0})\overline{F_{+}(\overline{z_{0}},z_{1})}dz_{0}\overline{dz_{1}}-\int\int_{\gamma}F_{-}(\overline{z_{1}},z_{0})F_{-}(\overline{z_{0}},z_{1})dz_{0}dz_{1}\right\}\end{array}\right$. (12)
Note that $c_{n}(z,z_{\partial})=c_{n}(z,z_{\partial},U)$ is a conformally invariant quantity, in the sense that if $f:U\to U^{\prime}$ is a conformal map, then $c_{n}(z,z_{\partial},U)=c_{n}(f(z),f(z_{\partial}),f(U))$. This follows from two facts: that each integrand is analytic (resp. antianalytic) in $z_{i}$ when $\sigma_{i}=1$ (resp. $\sigma_{i}=-1$), and that the functions $F_{\pm}=F_{\pm,U}$ satisfy $F_{\pm,U}(u,v)=(f^{\prime}(u))^{(\pm)}F_{\pm,f(U)}(f(u),f(v))$, where the superscript $(\pm)$ again denotes whether we take the complex conjugate or not. See Proposition 15 of [14], as well as the discussion after Proposition 20 of the same paper.
Note also that $c_{n}(z,z_{\partial})$ is always real, since the terms in the sum corresponding to $\sigma$ and $\sigma$ are complex conjugates. Moreover, since $F_{+}$ and $F_{-}$ are analytic on $U$ and remain bounded near $z_{\partial}$, the $c_{n}(z,z_{\partial})$ grow at most exponentially in $n$: there exists an absolute constant $C$ (depending only on $\gamma$ and $U$) such that
$\forall n\in\mathbb{Z}_{\geq 0},~c_{n}(z,z_{\partial})\leq C^{n}.\\$ (13)
@@
$\mathbbm{E}_{\varepsilon}\left[\binom{\frac{n_{\varepsilon}(z)-o_{\varepsilon}(z)}{2}}{n}\right]\geq\mathbbm{E}_{\varepsilon}\left[\binom{\frac{n_{\varepsilon}(z)-1}{2}}{n}\right$] $\geq\mathbbm{E}_{\varepsilon}\left[\mathds{1}_{n_{\varepsilon}(z)\geq 2n}\frac{(n_{\varepsilon}(z)-2n)^{n}}{2^{n}n!}\right]$
$\geq\mathbbm{E}_{\varepsilon}\left[\mathds{1}_{n_{\varepsilon}(z)\geq 4n}\frac{n_{\varepsilon}(z)^{n}}{4^{n}n!}\right]\geq\frac{1}{4^{n}n!}(\mathbbm{E}_{\varepsilon}[n_{\varepsilon}(z)^{n}]-(4n)^{n}).$
@@
since there are at most $n$ choices for $j_{1}$, $n/2$ choices for $j_{2}$ etc. Using Stierling’s formula, we conclude that $m_{n}\leq C^{n}n^{n}$ for another absolute constant $C$, hence Carleman’s condition is verified and $n_{\varepsilon}(z)$ converges towards a random variable $n(z)$ which is uniquely determined by its moments $m_{n}$. Since these moments are conformally invariant (they are expressed as integrals of conformal quantities), the law of $n(z)$ is also conformally invariant. ∎
In the particular case $U^{r}=[-\infty,+\infty]\times[-\pi/2,\pi/2]$ where the Green function has a simple explicit expression, the computations turn out to be particularly simple.
@@
The infinite strip does not fill the technical hypothesis of our theorem since it is not a bounded open set, although it fills the other conditions: it is symmetric and has a horizontal boundary. Nonetheless, since the formula for the moments depends only on conformally invariant quantities, the moments can be computed in any simply connected open set, and the computations transferred to any other domain by a Riemann map. Besides, if we take $U_{\varepsilon}^{r}=\varepsilon\mathcal{Z}^{2}\cap([-1/\varepsilon,1/\varepsilon]\times[-\pi/2,\pi/2])$, the proof of Theorem 3.4 still holds, at least for the folded dimer model: for the proof to extend to this folded case on $U_{\varepsilon}^{r}$, the only part of the proof that has to be modified is Lemma B.3 of the Appendix. The proof of this lemma still works because, to show that $K^{-1}(b,w)$ is close to the partial derivative of the Green function $g$, we only need that $g_{\mathbb{R}^{2}}$ (the full-plane Green function) or $g_{\mathbb{H}}$ (the half-plane Green function) and $g$ are $O(\varepsilon)$ on the boundary of $\mathcal{G}_{\varepsilon}^{r}$ to apply the Harnack lemma. This is also true here.
@@
$c_{1}(z,z_{\partial})=-4\Im\left\{\int_{x+iy}^{x+i\pi/2}\frac{1}{4\pi}dz\right\}=-\frac{1}{2}+\frac{y}{\pi}$